OpenStudy (mariebrittany07):

An employee was hired at a wage of $8.00 per hour. After a raise, the employee earned $8.75 cents per hour. What was the percent increase?

hero (hero):

If you let: a = 1st number b = 2nd number Then, the percent change from a to b is: \[\% \Delta = \frac{|a - b|}{a} \times 100\]

hero (hero):

There's nothing more to explain. Simply input a = 8.00 and b = 8.75 into the formula and evaluate.

OpenStudy (anonymous):

I'm looking at @Hero 's explanation and it is perfect. I don't think I could add anything to it.

OpenStudy (mariebrittany07):

can you show me more how to do it please

hero (hero):

\[\% \Delta = \frac{|8 - 8.75|}{8} \times 100\]

OpenStudy (mariebrittany07):

okay how do you solve it and evaluate it

OpenStudy (anonymous):

Are you familiar with the | | symbol for absolute value?

OpenStudy (mariebrittany07):

sort of

OpenStudy (anonymous):

A way-simplified explanation (which will suffice for now) is just to perform the operation(s) inside and make it positive, whatever the answer is. Then divide by 8, then multiply by 100 and you have a positive answer expressed as a percentage change.

OpenStudy (mariebrittany07):

so can you please show me

OpenStudy (anonymous):

What is 8 - 8.75 ?

OpenStudy (mariebrittany07):

\[8-8.75=-0.75\]

OpenStudy (mariebrittany07):

so then do i divide 0.75 by 8

hero (hero):

|n| = absolute value of a number = distance of a number from zero |-0.75| = absolute value of -0.75 = distance of -0.75 from zero = 0.75

OpenStudy (anonymous):

Ok, good. Yes, divide 0.75 by 8. Then multiply it by 100.

hero (hero):

Multiply by 100 first, then divide by 8

hero (hero):

You should have 75/8 afterwards

OpenStudy (mariebrittany07):

\[0.75 \div 8 = 0.09375 \]

OpenStudy (mariebrittany07):

\[0.09375\times 100 =937500\]

hero (hero):

Moving over two decimal places to the right is the same as multiplying by 100

OpenStudy (mariebrittany07):

am i right

hero (hero):

You should have 9.375

hero (hero):

How did the decimal point disappear after you multiplied by 100?

OpenStudy (mariebrittany07):

i have no idea how the decimal has disappeared after i multiplied it by 100

OpenStudy (mariebrittany07):

how did you get 9.375

hero (hero):

Try multiplying by 100 again

hero (hero):

Start with 0.09375, then multiply by 100

OpenStudy (mariebrittany07):

i have no idea what i did to get that answer but i got the right one now

OpenStudy (mariebrittany07):

is that all that i have to do

hero (hero):

Yes, you are essentially done, however, it is usually appropriate to state the percent change from a to b in the form of a sentence.

hero (hero):

Thus the percent change from a to b is 9.375